# Stochastic Online Conformal Prediction with Semi-Bandit Feedback

### v2

@@ 3 Algorithm
4 Experiments
4.1 Experimental Setup
Image classification task. We use the Vision Transformer [Dosovitskiy et al., 2020] model on the ImageNet dataset [Deng et al., 2009]. [^1] Obtained from https://www.image-net.org/ with a custom and non-commercial license; we use the $16\times 16$ down sampled version. Each image from the dataset belongs to exactly one class out of the 1,000 candidate classes. Consequently, the cardinality of the label domain is 1,000 (i.e., $|\mathcal{Y}|=1000$). In the experiments, images arrive sequentially, and our algorithm aims to construct the smallest candidate label set that achieves the desired coverage. We use the logits score returned by a ViT model (pretrained on the ImageNet training set) as our scoring function $f$; we use the logits instead of the softmax function to evaluate sensitivity to the choice of scoring function.
Document retrieval task. Next, we consider the Dense Passage Retriever (DPR) model [Karpukhin et al., 2020] on the SQuAD question-answering dataset. DPR leverages a dual-encoder architecture that maps questions and candidate documents to embedding vectors. Denoting the space of questions and documents by $\mathcal{Q}$ and $\mathcal{D}$, respectively, then DPR consists of a question encoder $E_{Q}:\mathcal{Q}\mapsto\mathbb{R}^{768}$ and a document encoder $E_{D}:\mathcal{D}\mapsto\mathbb{R}^{768}$. Given a question $q$ and a set of candidate documents $D\subseteq\mathcal{D}$, the similarity score between each document $d\in D$ and the question $q$ is defined as:
$s_{q,d}=\frac{{E_{Q}(q)\cdot E_{D}(d)}}{|E_{Q}(q)|\cdot|E_{D}(d)|}.$
Our dataset is SQuAD question-answering dataset [Rajpurkar et al., 2016], which is a popular benchmark for reading comprehension. Each question in SQuAD can be answered by finding the relevant information in a corresponding Wikipedia paragraph known as the context. The authors of DPR make a few additional changes to adapt SQuAD better for document retrieval. First, paragraphs are further split into multiple, disjoint text blocks of 100 words, serving as the basic retrieval unit (i.e., candidate documents). Second, each question is paired with ground truth documents and a set of irrelevant documents. [^2] In the data, all questions are paired with 50 irrelevant documents. We construct the candidate documents set by including all 50 irrelevant documents and 1 ground truth document. We exclude questions that have 0 ground truth documents. The data were obtained from DPR’s public Github Repo: https://github.com/facebookresearch/DPR with licenses CC BY-SA 4.0 and CC BY-NC 4.0 In our experiments, for each question, we include one ground truth document and all the irrelevant documents to create the set of candidate documents.
Second-price auctions task. Lastly, we consider the scenario of setting reservation prices in second-price auctions, a well-studied problem that has semi-bandit feedback [Cesa-Bianchi et al., 2014, Zhao and Chen, 2020]. In this problem, a seller (the auctioneer) repeatedly sells the same type of items to a group of bidders. In each round $t$, she publicly announces a reservation price $p^{t}$, while bidders draw their private values $v^{t}$ from a fixed distribution that is unknown to the seller. For each bidder $i$, she will submit the bid $B^{t}_{i}=v^{t}_{i}$ if and only if $v^{t}_{i}\geq p^{t}$. The seller obtains reward $R_{t}$:
That is, the seller aims to find the highest reservation price $p$ such that she can sell the item with probability at least $\alpha$. This problem can be solved using online conformal prediction. Following standard practice [Mohri and Medina, 2014], we use a synthetic dataset that adapts from eBay auction data [Jank and Shmueli, 2010]. Specifically, we simulate the distribution of $v_{i}$ by using the empirical distribution of the observed bids in the dataset.
Baselines. We compare to greedy, Adaptive Conformal Inference (ACI) [Gibbs and Candes, 2021], and Decaying Learning Rate (DLR) [Angelopoulos et al., 2024b]. The greedy strategy chooses
$\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{t}(\tau)\leq 1-\alpha\}$
at every step, which cannot guarantee the $\alpha$ coverage rate, leading it to undercover significantly more than desired. Next, Adaptive Conformal Inference (ACI) adjusts $\alpha_{t}$ (and then $\tau_{t}$) based on whether the ground truth label is in the previous round’s prediction set. We choose the learning rate $\gamma$ from a grid search in a candidate set proposed in [Gibbs and Candès, 2022]. To run ACI in our semi-bandit feedback setting, we only update the quantile function $\hat{Q}_{t}$ when ground-truth label $y_{t}^{*}$ is observed; as we show, this biased strategy for updating $\hat{Q}_{t}$ leads it to fail to achieve the desired coverage rate.
Lastly, we consider the Decaying Learning Rate (DLR) algorithm proposed in [Angelopoulos et al., 2024b]. In contrast to ACI, DLR directly performs gradient descent on the cutoffs $\tau_{t}$ instead of the quantiles $\alpha_{t}$; it can be viewed as running ACI with $\hat{Q}_{t}(\alpha)=\alpha$. We set the learning rate to the one used in the experiments from the original paper—i.e., $\eta_{t}=t^{-1/2-\epsilon}$ with $\epsilon=0.1$. We show results for two additional baselines, explore-then-commit (ETC) and conservative ETC, in Appendix B.
Experiment parameters. We use $\alpha=0.9$ and $T=10000$, and report the averages across 5 runs.
$\phi(\tau)=\begin{cases}-\lambda_{1}\cdot\lvert G^{*}(\tau)-(1-\alpha)\rvert&\text{if $G^{*}(\tau)\leq(1-\alpha)$ }\\ -\lambda_{2}\cdot\lvert G^{*}(\tau)-(1-\alpha)\rvert&\text{if $G^{*}(\tau)>(1-\alpha)$},\end{cases}$
@@ 4.2 Results
Regret. First, Figure 1 shows the cumulative regret of each approach on each task. As can be seen, our algorithm consistently obtains the lowest regret. The ACI algorithm attains a regret level comparable to the greedy algorithm because its quantile function is updated with a bias. Furthermore, note that the curves for both ACI and the greedy algorithm appear to be superlinear. This can happen since these algorithms do not properly account for semi-bandit feedback—in particular, the empirical estimate $G_{t}$ of the distribution becomes increasingly truncated.
Moreover, the poor performance of DLR can be attributed to the fact that it does not make use of a quantile function. As a consequence, without score-specific hyperparameter tuning (i.e., tuning the learning rate), it can converge very slowly to the optimal prediction sets. For many scoring functions, we do not have prior knowledge of the score’s range, which exacerbates these issues. These issues are particularly salient when we consider tasks such as the second-price auction, where the score’s range (i.e., the range of bids) can be difficult to predict in advance. In contrast, our algorithm consistently performs and does not have any hyperparamters to tune.
@@
Summary. These results demonstrate that our algorithm achieves (and converges to) the desired coverage rate while achieving sublinear regret and maintaining $\tau_{t}<\tau^{*}$. In contrast, ACI and greedy fail to achieve the desired coverage rate, and DLR converges much more slowly than our algorithm.
@@ 5 Conclusion
We have proposed a novel conformal prediction algorithm for constructing online prediction sets under stochastic semi-bandit feedback. We have shown our our algorithm can be applied to learn optimal prediction sets in image classification, document retrieval, and second-price auction reservation price prediction. Our experiments demonstrate that we maintain the desired $\alpha$ coverage level while achieving prediction set sizes that achieve sublinear regret and zero undercoverage count.
Appendix A Proof of Theorem 3.1
We begin by proving several helper lemmas, and then establish our main result. Our first lemma shows that $G_{t}$ defined in (3) converges uniformly to $G_{t}^{*}$ defined in (2). Note that the randomness comes from the sequence of random variables $s_{j}=f(x_{j},y_{j}^{*})$.
Lemma A.1.
For any $\epsilon\in\mathbb{R}_{>0}$, we have
$\mathbb{P}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right]\leq 2e^{-2t\epsilon}.$
Proof.
Let the full empirical CDF of $s$ be $G^{\prime}_{t}(\tau)=\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}(s_{j}\leq\tau)$, and define the events
$A_{<}$ =$=\left\{\sup_{\tau<\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}\qquad\text{and}\qquad A_{\geq}=\left\{\sup_{\tau\geq\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}.$
First, consider $A_{\geq}$. For $\tau\geq\tau_{t}$, by definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}^{*}(\tau)=G^{*}(\tau)$ and $G_{t}(\tau)=G^{\prime}_{t}(\tau)$. Thus, by the DKW inequality, we have $\mathbb{P}[A_{\geq}]\leq 2e^{-2t\epsilon^{2}}$. Next, consider $A_{<}$. By definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}(\tau)=G_{t}^{*}(\tau)=0$, so $\mathbb{P}[A_{<}]=0$. Thus, by a union bound, we have
$\mathbb{P}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right]$ $\leq\mathbb{P}[A_{\geq}]+\mathbb{P}[A_{<}]\leq 2e^{-2t\epsilon^{2}},$
as claimed. ∎
Lemma A.2.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, $\tau_{t}\leq\tau^{*}$ for all $t\in[T]$.
Proof.
We prove by induction. For the base case $t=1$, we choose $\tau_{1}=-\infty$, so $\tau_{1}\leq\tau^{*}$. For the inductive case, assume that $\tau_{k}\leq\tau^{*}$. Now, by our assumption that $\sup_{\tau\in\mathbb{R}}|G_{k}(\tau)-G_{k}^{*}(\tau)|\leq\epsilon_{k}$ for all $k\in[T]$, we have
$G^{*}_{k-1}(\tau)\leq G_{k-1}(\tau)+\epsilon_{k-1}=\overline{G}_{k-1}(\tau).$
Because $\tau_{k}\leq\tau^{*}$, we also have $G^{*}_{k-1}(\tau)=G^{*}(\tau)$ for all $\tau\geq\tau^{*}$. Together, we have
$\overline{G}_{k-1}(\tau^{*})\geq G^{*}(\tau^{*})=1-\alpha.$
Thus, we have $\tau_{1-\alpha,k}\leq\tau^{*}$. Since $\tau_{k}\leq\tau^{*}$, it follows that:
$\tau_{k+1}=\max\{\tau_{k-1},\tau_{1-\alpha,k}\}\leq\tau^{*},$
as claimed. ∎
Our next lemma bounds the range of $\overline{G}_{t}(\tau_{t})$.
Lemma A.3.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, for all $t\in[T]$, we have
Proof.
By Lemma A.2, $\tau_{t}\leq\tau^{*}$, implying that $G^{*}(\tau_{t})\leq G^{*}(\tau^{*})$. For the upper bound, we have
$\overline{G}_{t}(\tau_{t})\leq G_{t}^{*}(\tau_{t})+2\epsilon_{t}=G^{*}(\tau_{t})+2\epsilon_{t}\leq G^{*}(\tau^{*})+2\epsilon_{t}=1-\alpha+2\epsilon_{t}.$
as claimed. ∎
See 3.1
Proof.
Define the good event
By a union bound and by Lemma A.1, we have
Now, we have
$R_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}E\right]\cdot\mathbb{P}[E]+\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}\neg E\right]\cdot\mathbb{P}[\neg E]$
$=\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]+\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
$\leq\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]+\sum_{t=1}^{T}\mathbb{E}\left[2\phi_{max}\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
The first inequality follows from Assumption 2.2. Using Lemma A.3, we can bound $X_{t}$ as follows:
$X_{t}$ $=\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]$
$=\mathbb{E}[\psi(G^{*}(\tau^{*}))-\psi(G^{*}({\tau_{t}}))\mid E]$
$\leq K\cdot\mathbb{E}[|G^{*}(\tau^{*})-G^{*}(\tau_{t})|\mid E]$
$=K\cdot\mathbb{E}\left[|1-\alpha-\overline{G}_{t}(\tau_{t})+\overline{G}_{t}(\tau_{t})-G^{*}(\tau_{t})|\mid E\right]$
$\leq K\cdot\mathbb{E}\left[|1-\alpha-\overline{G}_{t}(\tau_{t})|+|\overline{G}_{t}(\tau_{t})-G^{*}(\tau_{t})|\mid E\right]$
$\leq K\cdot\mathbb{E}\left[\max\left\{\frac{2}{t},2\epsilon_{t}\right\}+|\overline{G}_{t}(\tau_{t})-G_{t}^{*}(\tau_{t})|\biggm{|}E\right]$
$\leq K\cdot\max\left\{\frac{2}{t},2\epsilon_{t}\right\}+K\cdot\mathbb{E}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\biggm{|}E\right]$
$\leq K\cdot\max\left\{\frac{2}{t},2\sqrt{\frac{\log(2/\delta)}{2t}}\right\}+2K\sqrt{\frac{\log(2/\delta)}{2t}}$
$\leq\frac{2K}{t}+4K\sqrt{\frac{\log(2/\delta)}{2t}},$
where the first inequality follows from Assumption 2.1 and the third inequality follows from Lemma A.3. Thus, we have
$\sum_{t=1}^{T}X_{t}$ $\leq\sum_{t=1}^{T}\left\{\frac{2K}{t}+4K\cdot\sqrt{\frac{\log T}{t}}\right\}\leq K\left[2\log T+1+4\sqrt{T\log T}\}\right].$
The claim follows. ∎
Appendix B Additional Experiments

### v3

@@ 3 Algorithm
$G_{t}(\tau)=\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}(\max\{\tau_{t},s_{j}\}\leq\tau).$ (3)
$\overline{G}_{t}(\tau)=G_{t}(\tau)+\epsilon_{t}.$ (4)
We use this $\overline{G}_{t}$ to compute $\tau_{t}$ in Algorithm 1.
Finally, Algorithm 1 satisfies the following theoretical guarantee: (i) it incurs sublinear regret $\tilde{\mathcal{O}}(\sqrt{T})$, and (ii) it satisfies $C_{t}^{*}\subseteq C_{t}$ for all $t$ with probability at least $1-2/T$: {restatable}theoremthmMain Algorithm 1 satisfies
$R_{T}\leq K\left(2\log T+4\sqrt{T\log T}+1\right)+4\phi_{max},$
Also, with probability at least $1-2/T$, $\tau_{t}\leq\tau^{*}$ for all $t$.
4 Proof of Theorem 1
Our first lemma shows that $G_{t}$ defined in (3) converges uniformly to $G_{t}^{*}$ defined in (2). Note that the randomness comes from the sequence of random variables $s_{j}=f(x_{j},y_{j}^{*})$.
Lemma 4.1.
For any $\epsilon\in\mathbb{R}_{>0}$, $\mathbb{P}[E_{t}]\leq 2e^{-2t\epsilon}$, where
$E_{t}=\left\{\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}$
Proof.
Let the empirical CDF of $s$ be $G^{\prime}_{t}(\tau)=\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}(s_{j}\leq\tau)$, and define the events
$A_{<}$ =$=\left\{\sup_{\tau<\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}$
$A_{\geq}$ =$=\left\{\sup_{\tau\geq\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}.$
First, consider $A_{\geq}$. For $\tau\geq\tau_{t}$, by definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}^{*}(\tau)=G^{*}(\tau)$ and $G_{t}(\tau)=G^{\prime}_{t}(\tau)$. Thus, by the DKW inequality, $\mathbb{P}[A_{\geq}]\leq 2e^{-2t\epsilon^{2}}$. Next, consider $A_{<}$. By definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}(\tau)=G_{t}^{*}(\tau)=0$, so $\mathbb{P}[A_{<}]=0$. Thus, by a union bound, we have
$\mathbb{P}[E_{t}]\leq\mathbb{P}[A_{\geq}]+\mathbb{P}[A_{<}]\leq 2e^{-2t\epsilon^{2}},$
as claimed. ∎
Lemma 4.2.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$ with $\epsilon_{t}=\sqrt{\log(2/\delta)/2t}$Then, $\tau_{t}\leq\tau^{*}$ for all $t\in[T]$.
Proof.
We prove by induction. For the base case, when $t=1$, we set $\tau_{1}=-\infty$. Thus, $\tau_{1}\leq\tau^{*}$. For the inductive case, note that $\tau_{k}\leq\tau^{*}$ by the inductive hypothesis, so we have $G_{k}^{*}(\tau)=G^{*}(\tau)$ for all $\tau\geq\tau^{*}$. Then, we have $|G_{k}(\tau^{*})-G^{*}(\tau^{*})|=|G_{k}(\tau^{*})-G_{k}^{*}(\tau^{*})|\leq\sup_{\tau\in\mathbb{R}}|G_{k}(\tau)-G_{k}^{*}(\tau)|\leq\epsilon_{k}$, where the last inequality follows from our assumption. Thus, we have $\overline{G}_{k}(\tau^{*})=G_{k}(\tau^{*})+\epsilon_{k}\geq G^{*}(\tau^{*})=1-\alpha$. Recall that we define $\tau_{1-\alpha,k}=\sup\{\tau\in\mathbb{R}\mid\overline{G}_{k}(\tau)\leq 1-\alpha\}$; thus, we have $\tau_{1-\alpha,k}\leq\tau^{*}$. Since $\tau_{k+1}=\max\{\tau_{1-\alpha,k},\tau_{k}\}$ with $\tau_{1-\alpha,k}\leq\tau^{*}$ and $\tau_{k}\leq\tau^{*}$, we have $\tau_{k+1}\leq\tau^{*}$. ∎
Our next lemma bounds the range of $\overline{G}_{t}(\tau_{t})$.
Lemma 4.3.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$ with $\epsilon_{t}=\sqrt{\log(2/\delta)/2t}$. Then, for all $t\in[T]$, we have
Proof.
By Lemma 4.2, $\tau_{t}\leq\tau^{*}$, implying that $G^{*}(\tau_{t})\leq G^{*}(\tau^{*})$. For the upper bound, we have
$\overline{G}_{t}(\tau_{t})\leq G_{t}^{*}(\tau_{t})+2\epsilon_{t}=G^{*}(\tau_{t})+2\epsilon_{t}$ $\leq G^{*}(\tau^{*})+2\epsilon_{t}$
$=1-\alpha+2\epsilon_{t}.$
as claimed. ∎
[…]
Proof.
Define the good event
By a union bound and by Lemma 4.1, we have
Now, we have
$R_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}E\right]\mathbb{P}[E]$
$\qquad+\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}\neg E\right]\mathbb{P}[\neg E]$
$=\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\mathbb{P}[E]$
$\qquad+\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid\neg E\right]\mathbb{P}[\neg E]$
$\leq\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\mathbb{P}[E]$
$\qquad+\sum_{t=1}^{T}\mathbb{E}\left[2\phi_{max}\mid\neg E\right]\mathbb{P}[\neg E]$
The first inequality follows from Assumption 2.2. Using Lemma 4.3, we can bound $X_{t}$ as follows:
$X_{t}$ $=\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]$
$=\mathbb{E}[\psi(G^{*}(\tau^{*}))-\psi(G^{*}({\tau_{t}}))\mid E]$
$\leq K\mathbb{E}[|G^{*}(\tau^{*})-G^{*}(\tau_{t})|\mid E]$
$=K\mathbb{E}\left[|1-\alpha-\overline{G}_{t}(\tau_{t})+\overline{G}_{t}(\tau_{t})-G^{*}(\tau_{t})|\mid E\right]$
$\leq K\mathbb{E}\left[|1-\alpha-\overline{G}_{t}(\tau_{t})|+|\overline{G}_{t}(\tau_{t})-G^{*}(\tau_{t})|\mid E\right]$
$\leq K\mathbb{E}\left[\max\left\{\frac{2}{t},2\epsilon_{t}\right\}+|\overline{G}_{t}(\tau_{t})-G_{t}^{*}(\tau_{t})|\biggm{|}E\right]$
$\leq K\max\left\{\frac{2}{t},2\epsilon_{t}\right\}+2K\mathbb{E}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\biggm{|}E\right]$
$\leq K\max\left\{\frac{2}{t},2\sqrt{\frac{\log(2/\delta)}{2t}}\right\}+2K\sqrt{\frac{\log(2/\delta)}{2t}}$
$\leq\frac{2K}{t}+4K\sqrt{\frac{\log(2/\delta)}{2t}},$
where the first inequality follows from Assumption 2.1 and the third inequality follows from Lemma 4.3. Thus, we have
$\sum_{t=1}^{T}X_{t}$ $\leq\sum_{t=1}^{T}\left\{\frac{2K}{t}+4K\sqrt{\frac{\log T}{t}}\right\}$
$\leq K\left[2\log T+1+4\sqrt{T\log T}\}\right].$
The claim follows. ∎
@@ 4 Proof of Theorem 1
5 Experiments
5.1 Experimental Setup
Image classification task. We use the Vision Transformer (Dosovitskiy et al., 2020) model on the ImageNet dataset (Deng et al., 2009). [^1] Obtained from https://www.image-net.org/ with a custom and non-commercial license; we use the $16\times 16$ down sampled version. Each image from the dataset belongs to exactly one class out of the 1,000 candidate classes. Consequently, the cardinality of the label domain is 1,000 (i.e., $|\mathcal{Y}|=1000$). In the experiments, images arrive sequentially, and our algorithm aims to construct the smallest candidate label set that achieves the desired coverage. We use the logits score returned by a ViT model (pretrained on the ImageNet training set) as our scoring function $f$; we use the logits instead of the softmax function to evaluate sensitivity to the choice of scoring function.
Document retrieval task. Next, we consider the Dense Passage Retriever (DPR) model (Karpukhin et al., 2020) on the SQuAD question-answering dataset. DPR leverages a dual-encoder architecture that maps questions and candidate documents to embedding vectors. Denoting the space of questions and documents by $\mathcal{Q}$ and $\mathcal{D}$, respectively, then DPR consists of a question encoder $E_{Q}:\mathcal{Q}\mapsto\mathbb{R}^{768}$ and a document encoder $E_{D}:\mathcal{D}\mapsto\mathbb{R}^{768}$. Given a question $q$ and a set of candidate documents $D\subseteq\mathcal{D}$, the similarity score between document $d\in D$ and question $q$ is
$s_{q,d}=\frac{{E_{Q}(q)E_{D}(d)}}{|E_{Q}(q)|\cdot|E_{D}(d)|}.$
Our dataset is SQuAD question-answering dataset (Rajpurkar et al., 2016), a popular reading comprehension benchmark. Each question in SQuAD can be answered by finding the relevant information in a corresponding Wikipedia paragraph known as the context. The authors of DPR make a few changes to adapt SQuAD to document retrieval. First, paragraphs are further split into multiple, disjoint text blocks of 100 words, serving as the basic retrieval unit (i.e., candidate documents). Second, each question is paired with ground truth documents and a set of irrelevant documents. [^2] In the data, all questions are paired with 50 irrelevant documents. We construct the candidate documents set by including all 50 irrelevant documents and 1 ground truth document. We exclude questions that have 0 ground truth documents. The data were obtained from DPR’s public Github Repo: https://github.com/facebookresearch/DPR with licenses CC BY-SA 4.0 and CC BY-NC 4.0 In our experiments, for each question, we include one ground truth document and all the irrelevant documents to create the set of candidate documents.
Second-price auctions task. Lastly, we consider the scenario of setting reservation prices in second-price auctions, a well-studied problem that has semi-bandit feedback (Cesa-Bianchi et al., 2014; Zhao & Chen, 2020). In this problem, a seller (the auctioneer) repeatedly sells the same type of items to a group of bidders. In each round $t$, she publicly announces a reservation price $p^{t}$, while bidders draw their private values $v^{t}$ from a fixed distribution that is unknown to the seller. For each bidder $i$, she will submit the bid $B^{t}_{i}=v^{t}_{i}$ if and only if $v^{t}_{i}\geq p^{t}$. The seller obtains reward $R_{t}$:
That is, the seller aims to find the highest reservation price $p$ such that she can sell the item with probability at least $\alpha$. This problem can be solved using online conformal prediction. Following standard practice (Mohri & Medina, 2014), we use a synthetic dataset that adapts from eBay auction data (Jank & Shmueli, 2010). Specifically, we simulate the distribution of $v_{i}$ by using the empirical distribution of the observed bids in the dataset.
Baselines. We compare to greedy, Adaptive Conformal Inference (ACI) (Gibbs & Candes, 2021), and Decaying Learning Rate (DLR) (Angelopoulos et al., 2024b). The greedy strategy chooses
$\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{t}(\tau)\leq 1-\alpha\}$
at every step, which cannot guarantee the $\alpha$ coverage rate, leading it to undercover significantly more than desired. Next, Adaptive Conformal Inference (ACI) adjusts $\alpha_{t}$ (and then $\tau_{t}$) based on whether the ground truth label is in the previous round’s prediction set. We choose the learning rate $\gamma$ from a grid search in a candidate set proposed in (Gibbs & Candès, 2024). To run ACI in our semi-bandit feedback setting, we only update the quantile function $\hat{Q}_{t}$ when ground-truth label $y_{t}^{*}$ is observed; as we show, this biased strategy for updating $\hat{Q}_{t}$ leads it to fail to achieve the desired coverage rate.
Lastly, we consider the Decaying Learning Rate (DLR) algorithm proposed in (Angelopoulos et al., 2024b). In contrast to ACI, DLR directly performs gradient descent on the cutoffs $\tau_{t}$ instead of the quantiles $\alpha_{t}$; it can be viewed as running ACI with $\hat{Q}_{t}(\alpha)=\alpha$. We set the learning rate to the one used in the experiments from the original paper—i.e., $\eta_{t}=t^{-1/2-\epsilon}$ with $\epsilon=0.1$. We show results for two additional baselines, explore-then-commit (ETC) and conservative ETC, in Appendix A.
Experiment parameters. We use $\alpha=0.9$ and $T=10000$, and report averages across 10 runs. The shaded areas are the 95% confidence intervals computed from 10 runs.
$\phi(\tau)=\begin{cases}-\lambda_{1}\lvert G^{*}(\tau)-(1-\alpha)\rvert&\text{if $G^{*}(\tau)\leq(1-\alpha)$ }\\ -\lambda_{2}\lvert G^{*}(\tau)-(1-\alpha)\rvert&\text{if $G^{*}(\tau)>(1-\alpha)$},\end{cases}$
@@ 5.2 Results
Regret. First, Figure 1 shows the cumulative regret of each approach on each task. We applied the log-scale for better readability. As can be seen, our algorithm consistently obtains the lowest regret. The ACI algorithm attains a regret level comparable to the greedy algorithm because its quantile function is updated with a bias. Furthermore, note that the curves for both ACI and the greedy algorithm appear to be superlinear. This can happen since these algorithms do not properly account for semi-bandit feedback—in particular, the empirical estimate $G_{t}$ of the distribution becomes increasingly truncated. Moreover, the poor performance of DLR can be attributed to the fact that it does not make use of a quantile function. As a consequence, without score-specific hyperparameter tuning (i.e., tuning the learning rate), it can converge very slowly to the optimal prediction sets. For many scoring functions, we do not have prior knowledge of the score’s range, which exacerbates these issues. These issues are particularly salient when we consider tasks such as the second-price auction, where the score’s range (i.e., the range of bids) can be difficult to predict in advance. In contrast, our algorithm consistently performs and does not have any hyperparamters to tune.
@@
Summary. These results show that our algorithm achieves (and converges to) the desired coverage rate while achieving sublinear regret and maintaining $\tau_{t}<\tau^{*}$. In contrast, ACI and greedy fail to achieve the desired coverage rate, and DLR converges much more slowly than our algorithm.
@@ 6 Conclusion
We have proposed a novel conformal prediction algorithm for constructing online prediction sets under stochastic semi-bandit feedback. We have shown our algorithm can be applied to learn optimal prediction sets in image classification, document retrieval, and second-price auction reservation price prediction. Our experiments show we achieve the desired $\alpha$ coverage level while achieving prediction set sizes that achieve sublinear regret and zero undercoverage count.
Impact Statement
Our approach aims to improve the trustworthiness of machine learning models via reliably uncertainty quantification; we do not foresee any ethical concerns.
Appendix A Additional Experiments